Consider two data sets. Set A: n = 5; x = 4 Set B: n = 50; x = 4 (a) Suppose the number 14 is included as an additional data value in Set A. Compute x for the new data set. Hint: x = nx. To compute x for the new data set, add 14 to x of the original data set and divide by 6. (Round your answer to two decimal places.) (b) Suppose the number 14 is included as an additional data value in Set B. Compute x for the new data set. (Round your answer to two decimal places.)
Answer #1

a) New mean for set A = (5*4 + 14)/6 = 36/6 = 6

b) New mean for set B = (50*4 + 14)/51 = 4.20